OpenStudy (anonymous):

Simplify the complex fractions by any method 4y+1/9y^4/y+2/y^3

OpenStudy (psymon):

\[\frac{ \frac{ 4y+1 }{ 9y ^{4} } }{ \frac{ y+2 }{ y ^{3} } }=\frac{ 4y+1 }{ 9y ^{4} }\div \frac{ y+2 }{ y ^{3} }\] If you know how to do a normal division of fractions, then maybe that makes more sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 4y+1 }{ 9y^4 }\div \frac{ y+2 }{y^3 }\]

OpenStudy (anonymous):

\[\frac{ 4y+1 }{ 9y^4 } \times \frac{ y^3 }{ y+2}\]

OpenStudy (psymon):

Yep, exactly. Did that help ya simplify it some?

OpenStudy (anonymous):

still trying

OpenStudy (psymon):

Alrighty, no worries :3

OpenStudy (anonymous):

what is :3

OpenStudy (psymon):

Oh, its a smiley face xD If you look at it sideways, lol.

OpenStudy (anonymous):

lol

OpenStudy (psymon):

Yeah, my bad. Just let me know if you want me to work through it more or something.

OpenStudy (anonymous):

will do